OpenStudy (anonymous):

By what number must you multiply each side of the equation 4/5x=8 to produce the equivalent equation x = 10? 5/4 5 1/5 -4/5

OpenStudy (skullpatrol):

Hi

OpenStudy (anonymous):

i don't understand this one may any of u help usually i have a answer but on this one pure confusion

OpenStudy (anonymous):

hey @skullpatrol

undeadknight26 (undeadknight26):

i has no idea...

OpenStudy (anonymous):

but if i had to guess it would be B

undeadknight26 (undeadknight26):

hmmm nope wait a second... 4/5x=8 /4 /4 5x = 2 *5 *5 x = 10...no its not that...i have no idea...im not sure about what i just did but im quite sure its not right...

OpenStudy (anonymous):

or C 1/5

OpenStudy (anonymous):

may anyone help me with this

OpenStudy (anonymous):

4/5 (1/5)x=8(1/5) 4/25x=8/5 4x=8/5(25) 4x=40 x=10 1/5 is correct

OpenStudy (skullpatrol):